OpenStudy (anonymous):

(18x^5 + 6x^4 − 12x^3) ÷ 6x^2

jimthompson5910 (jim_thompson5910):

\[\large (18x^5 + 6x^4 - 12x^3) \div 6x^2\] \[\large \frac{18x^5 + 6x^4 - 12x^3}{6x^2}\] \[\large \frac{6x^2(3x^3 + x^2 - 2x)}{6x^2}\] I'll let you finish up

OpenStudy (anonymous):

I have no idea how to do this @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Ok here's how you do the full problem \[\large (18x^5 + 6x^4 - 12x^3) \div 6x^2\] \[\large \frac{18x^5 + 6x^4 - 12x^3}{6x^2}\] \[\large \frac{6x^2(3x^3 + x^2 - 2x)}{6x^2}\] \[\large \frac{\cancel{6x^2}(3x^3 + x^2 - 2x)}{\cancel{6x^2}}\] \[\large \frac{1(3x^3 + x^2 - 2x)}{1}\] \[\large \frac{3x^3 + x^2 - 2x}{1}\] \[\large 3x^3 + x^2 - 2x\]

jimthompson5910 (jim_thompson5910):

hopefully that helps

jimthompson5910 (jim_thompson5910):

the idea is to factor out the 6x^2 since that's in the denominator then you cancel out the common terms and simplify

OpenStudy (anonymous):

Thanks so much

jimthompson5910 (jim_thompson5910):

np

OpenStudy (lncognlto):

Can it be maybe be simplified further, though?

OpenStudy (lncognlto):

As in 3x^3 + x^2 - 2x = x(3x^2 + x - 2) = x(3x - 2)(x+1)

jimthompson5910 (jim_thompson5910):

you could factor, but it's pretty much simplified at this point

OpenStudy (lncognlto):

Yes, sorry, I meant maybe it could be factored.